Determine the x and y intercept of the line whose equation is 2x-3y- 5/4 =0.
1
Expert's answer
2020-09-24T15:07:03-0400

"2x-3y- 5\/4 =0"

"y=\dfrac{8x-5}{12}"

The y- intercept of the line in point "A(0,y_1)"

Consider "A(0;y_1)" in "y=\dfrac{8x-5}{12}"


"y_1=\dfrac{8*0-5}{12}=-\dfrac{5}{12}"

"A(0,-\dfrac{5}{12})"


The x- intercept of the line in point "B(x_1,0)"

Consider "B(x_1,0)" in "y=\dfrac{8x-5}{12}"


"0=\dfrac{x_1*8-5}{12}"

"x_1*8-5=0"

"x_1=\dfrac{5}{8}"

"B(\dfrac{5}{8},0)"